Solve the initial value problem below using the method of Laplace transforms. y - 12y +45y = 39 e 4t, y(0) = 3, y(0) = 15
Answer #1

Quee ylizy +484 = 3ge at 4(0)- ay(o)=15 dyt sees)-yeo) mering R$ } o s2 Yee) -s Yeo) -y.com - 5-4 Take saplace transformA Bro, Ca 3, no YOS) 2 3 S-4 + 2 125 +45 - Set 125 +36-36405 Yn = 3 + cajet 52 3 + 3 (6-6)2 32 S-4 ylt) = 3e4 on37 ) 3 e 4th

If you have any questions please let me know

Thank you

Please give me thumb up..